the longest wavelenth of He^++ in Paschen series is m the shortest wavelegth of Be^3+ in paschen series (in terms of m)

Grade:11

1 Answers

Mohammed Imran
6 years ago
As 1/wavelength=Rz^2(1/n1^2-1/n2^2)Longest wavelength of paschen series in HE^3+ n1=3 , n2=4 , R = Rydberg contact, z- atomic number as z of HE = 2 Put it in formula you will get wavelength m=36/7R Now for shortest wavelength of BE^3+ in paschen series n1= 3 , n2=infinity ,Z of beryllium is 4Put it in formula you will get wavelength = 9/16R Final answer is wavelength of BE^3+ in terms of m = 7m/64 Final wavelength = 7m/64